The following table shows the value of credits and debits in four balance-of-payments accounts in the fictional economy of Gamros.

Balance-of-Payments Account

Debits

Credits

(Billions of dollars)

(Billions of dollars)

Merchandise 60 40
Services 25 10
Income Receipts 45 30
Unilateral Transfers 15 15
Total 145 95

Based on the total values of Gamros’s debits and credits, there is   money flowing out of Gamros to foreign countries than there is flowing into Gamros. This means that Gamros has a deficit in the   account.

Complete the following table based on the credits and debits presented in the previous table.

Term

How to Calculate

Deficit or Surplus

Size of Deficit or Surplus

(Billions of dollars)

Balance of Trade      
Net Exports      
Answer #1

The Balance of Trade or net exports is the difference between the value of a country's exports of goods and services and the country's imports for the same. In this example, we have exports as Merchandise + Service = 40 + 10 = 50. Imports are = 60 + 25 = 85

THUS, Balance of Trade = 50 - 85 = -35

The country is running a balance of trade deficit i .e. it is importing more than it is exporting